# Tool-Aided Evolutionary LLM for Generative Policy Toward Efficient Resource Management in Wireless Federated Learning

### v1

@@ VI-B Performance Comparison
We first show the overall performance comparison, in terms of CWEPM in (13a), learning accuracy, and energy consumption per round, and Fig. 5 presents the corresponding results. Notably, as shown in Fig. 6, significantly heterogeneous label distributions and dataset sizes exist for evaluation. As shown in Fig. 5(a), the proposed T-ELLM GRPO and T-ELLM PPO achieve the highest CWPEM values, suggesting that these methods require less energy to attain target performance levels. Fig. 5(b) illustrates the convergence of wireless FL under different algorithms. The proposed T-ELLM GRPO and T-ELLM PPO exhibit a remarkable convergence rate towards the desired accuracy. Additionally, Fig. 5(c) presents the energy consumption per round of wireless FL, where the solid curves therein represent the mean testing accuracy across ten experimental trials and the shaded part represents the $95\%$ confidence interval calculated from these experiments. The proposed T-ELLM GRPO and T-ELLM PPO consume less energy per round, demonstrating their efficiency in resource utilization during the FL process. Furthermore, the narrower confidence interval observed for T-ELLM GRPO and T-ELLM PPO indicates more stable and consistent decision-making.
@@ VI-C1 Adaptation to Resource Changes
To verify the generalization of the proposed T-ELLM to the environment, we first change the computing and communication capabilities of each device in the environment. Notably, Fig. 8 highlights the differences in terms of the maximum CPU frequency and transmission power of devices. Compared to the environment (i.e., Scenario 1) for evaluation in Section VI-B, the modified environment is termed Scenario 2. Notably, all algorithms are directly transferred from the training outcome in Scenario 1 and have no prior interaction with Scenario 2, ensuring an unbiased assessment of adaptability. The corresponding performance in Scenario 2 is shown in Fig. 9. As demonstrated in Fig. 9(a), the CWPEM of T-ELLM GRPO and T-ELLM PPO outperforms that of other baseline algorithms, indicating superior performance with lower energy consumption. Furthermore, Fig. 9(b) confirms that the learning accuracy of T-ELLM remains robust, without sacrificing the learning efficiency. Furthermore, Fig. 9(c) shows that the energy consumption per round yielded by T-ELLM is also lower than that of other algorithms, demonstrating its ability to maintain high accuracy while optimizing energy efficiency. These results highlight T-ELLM’s adaptability to varying environmental conditions. Additionally, the FedAvg tool and greedy are not learning-based algorithms, naturally, their performance is less affected by environmental changes. In comparison, as shown in Fig. 9(c), due to its training only in the default Scenario 1, the energy consumption of FL-DLT3 is significantly higher than that of other algorithms. This further underscores the advantages of the proposed T-ELLM in terms of environmental adaptability and energy efficiency.
@@ VI-C2 Adaptation to Changed Task Requirements
In this part, we further demonstrate the ability of T-ELLM to cope with different QoS requirements. Fig. 10 provides the results after changing $T_{\text{QoS}}$ in the default Scenario 1 from $15$ seconds to $20$ seconds. The results show that the proposed T-ELLM GRPO and T-ELLM PPO achieve the highest CWPEM values in Fig. 10(a), indicating superior energy efficiency in wireless FL. Fig. 10(b) illustrates the convergence of wireless FL under different algorithms. The proposed T-ELLM GRPO and T-ELLM PPO exhibit a significant convergence rate, achieving the desired accuracy. Fig. 10(c) shows that the proposed T-ELLM GRPO and T-ELLM PPO consume less energy per round, while it can be seen that the FedAvg Tool also maintains a low energy consumption per round. The reason lies in that the resource tool is used therein for resource allocation, contributing to saving overall energy consumption required for the anticipated QoS. On the other hand, although the $\epsilon$-Greedy tool algorithm also gets help from the tool, the extensive reliance on the artificial setting of greedy strategy parameters undermines the potential benefit. Additionally, as shown in Fig. 10(c), FL-DLT3 demonstrates notably higher energy consumption, as it depends solely on its NNs for resource allocation and lacks adaptability to varying QoS demands. In comparison, the proposed T-ELLM framework dynamically adjusts to QoS requirements, enabling more efficient decision-making.
@@
We also evaluate the performance of the proposed T-ELLM with different QoS requirements and heterogeneity. The performance of the proposed T-ELLM is shown in Fig. 11. It can be seen in Fig. 11(a) that the proposed T-ELLM maintains comparable accuracy with different QoS requirements. Furthermore, the energy consumption decreases with the increase of QoS requirements. Fig. 11(b) shows the performance of the proposed T-ELLM with different data heterogeneity, and similar observations can be attained. These results show that the proposed T-ELLM can adapt to environmental changes to yield satisfactory accuracy in wireless FL.
VII Conclusion
@@
We have presented T-ELLM, a novel tool-aided evolutionary LLM framework for efficient device selection and resource allocation in wireless FL. Notably, T-ELLM has been developed on top of a mathematics-driven decoupling of LLM-based device selection and tool-based resource allocation. Meanwhile, the combination of the linguistic reasoning capabilities in LLMs and mathematical optimization tools contributes to boosting the generalization capability of decision-making in environmental changes. In addition, T-ELLM takes advantage of a model-based virtual environment to support GRPO-based fine-tuning at minimal communication cost during interactions. Our theoretical analysis has proved the bounded discrepancy between virtual and real environments, which ensures the advantage function learned in the virtual environment maintains a provably small deviation from real-world conditions. Extensive experimental results have demonstrated that T-ELLM can further improve the energy efficiency and exhibit robust adaptability to environmental changes.

### v2

@@ VI-B Performance Comparison
We first show the overall performance comparison, in terms of CWEPM in (8a), learning accuracy, and energy consumption per round, and Fig. 5 presents the corresponding results. Notably, as shown in Fig. 6, significantly heterogeneous label distributions and dataset sizes exist for evaluation. As shown in Fig. 5(a), the proposed T-ELLM GRPO and T-ELLM PPO achieve the highest CWPEM values, suggesting that these methods require less energy to attain target performance levels. Fig. 5(b) illustrates the convergence of wireless FL under different algorithms. The proposed T-ELLM GRPO and T-ELLM PPO exhibit a remarkable convergence rate towards the desired accuracy. Additionally, Fig. 5(c) presents the energy consumption per round of wireless FL, where the solid curves therein represent the mean testing accuracy across $10$ experimental trials and the shaded part represents the $95\%$ confidence interval calculated from these experiments (the other figures are likewise). The proposed T-ELLM GRPO and T-ELLM PPO consume less energy per round, demonstrating their efficiency in resource utilization during the FL process. Furthermore, the narrower confidence interval observed for T-ELLM GRPO and T-ELLM PPO, relative to the FL-DLT3 and Fedavg Tool baseline, indicates stable and consistent decision-making, while maintaining a level of stability comparable to the heuristic methods.
@@ VI-C1 Adaptation to Resource Changes
To verify the generalization of the proposed T-ELLM to the environment, we first change the computing and communication capabilities of each device in the environment. Fig. 8 highlights the differences in terms of the maximum CPU frequency and transmission power of devices. Compared to the environment (i.e., Scenario 1) for evaluation in Section VI-B, the modified environment is termed Scenario 2. Notably, all algorithms are directly transferred from the training outcome in Scenario 1 and have no prior interaction with Scenario 2, ensuring an unbiased assessment of adaptability. The corresponding performance in Scenario 2 is shown in Fig. 9. As demonstrated in Fig. 9(a), the CWPEM of T-ELLM GRPO and T-ELLM PPO outperforms that of other baseline algorithms, indicating superior performance with lower energy consumption. Furthermore, Fig. 9(b) confirms that the learning accuracy of T-ELLM remains robust, without sacrificing the learning efficiency. Furthermore, Fig. 9(c) shows that the energy consumption per round yielded by T-ELLM is also lower than that of other algorithms, demonstrating its ability to maintain high accuracy while optimizing energy efficiency. These results highlight T-ELLM’s adaptability to varying environmental conditions. Additionally, the FedAvg tool and $\epsilon$-Greedy tool are not learning-based algorithms, naturally, their performance is less affected by environmental changes. In comparison, as shown in Fig. 9(c), due to its training only in the default Scenario 1, the energy consumption of FL-DLT3 and SAC is significantly higher than that of other algorithms. This further underscores the advantages of the proposed T-ELLM in terms of environmental adaptability and energy efficiency.
@@ VI-C2 Adaptation to Changed Task Requirements
In this part, we further demonstrate the ability of T-ELLM to cope with different QoS requirements. Fig. 10 provides the results after changing $T_{\text{QoS}}$ in the default Scenario 1 from $15$ seconds to $20$ seconds. The results show that the proposed T-ELLM GRPO and T-ELLM PPO achieve the highest CWPEM values in Fig. 10(a), indicating superior energy efficiency in wireless FL. Fig. 10(b) illustrates the convergence of wireless FL under different algorithms. The proposed T-ELLM GRPO and T-ELLM PPO exhibit a significant convergence rate, achieving the desired accuracy. Fig. 10(c) shows that the proposed T-ELLM GRPO and T-ELLM PPO consume less energy per round, while it can be seen that the FedAvg Tool also maintains a low energy consumption per round. The reason lies in that the resource tool is used therein for resource allocation, contributing to saving overall energy consumption required for the anticipated QoS. On the other hand, although the $\epsilon$-Greedy tool algorithm also gets help from the tool, the extensive reliance on the artificial setting of greedy strategy parameters undermines the potential benefit. Additionally, as shown in Fig. 10(c), FL-DLT3 and SAC demonstrate notably higher energy consumption, as they depend solely on their NNs for resource allocation and lack adaptability to varying QoS demands. In comparison, the proposed T-ELLM framework dynamically adjusts to QoS requirements, enabling more efficient decision-making.
@@
We also evaluate the performance of the proposed T-ELLM with different QoS requirements and heterogeneity. The performance of the proposed T-ELLM is shown in Fig. 11. It can be seen in Fig. 11(a) that the proposed T-ELLM maintains comparable accuracy with different QoS requirements. Furthermore, the energy consumption decreases with the increase of QoS requirements. Fig. 11(b) shows the performance of the proposed T-ELLM with different levels of data heterogeneity, and similar observations can be attained. These results show that the proposed T-ELLM can adapt to environmental changes by yielding appealing accuracy.
VI-D Performance of TELLM in Different Model Sizes
(a) Energy efficiency
(b) Convergence of Wireless FL
(c) Energy consumption per round
Figure 12: Performance comparison of large models with different scales for wireless FL.
In order to validate the effect of LLM size and language training approach on the T-ELLM, we evaluate the performance of three versions of GPT-2, including GPT-2 Small, GPT-2 Medium, and GPT-2 Large, corresponding to parameter scales of 0.10.1B, 0.30.3B, and 0.70.7B, respectively. Additionally, we compare the performance of GPT-2 models against a more lightweight, task-specific baseline we call GPT-Linear, which uses a Transformer decoder-only architecture (like GPT-2). The result is shown in Fig. 12. It can be seen that the performance for the GPT-2 series is all very similar, with no significant difference. Compared with the GPT-2 series, the CWPEM and convergence results of GPT-Linear are slightly lower, while the confidence interval of GPT-Linear is narrower than that of the GPT-2 series. These results validate the advantage of adopting a linguistic model, and a computationally efficient GPT-2 Small model already yields satisfactory results.
VI-E The Adaptability of TELLM to Different FL Algorithms
We investigate the adaptability and robustness of our T-ELLM framework for FL with gradient quantization[10] and Differential Privacy (DP)[4].
VI-E1 Adaptation to Gradient Quantization
We present the performance of T-ELLM with varying levels of gradient quantization in Fig. 13. As shown in Fig. 13(a), consistent with our intuition, the energy efficiency of T-ELLM decreases as the number of quantization bits decreases, which is expected since lower bit quantization introduces more noise, potentially degrading model accuracy. However, even with $2$-bit quantization, T-ELLM maintains a reasonable balance between energy consumption and accuracy. Fig. 13(b) shows that T-ELLM converges to a satisfactory accuracy level across all quantization levels, although higher bit quantization leads to faster convergence and higher final accuracy. Fig. 13(c) indicates that energy consumption per round increases with the number of quantization bits, as higher precision requires more data transmission. Overall, T-ELLM demonstrates robust performance even under aggressive gradient quantization.
VI-E2 Adaptation to Differential Privacy
We implement $(\epsilon,\delta)$-DP in the wireless FL process, where $\delta$ accounts for the probability that plain $\epsilon$-DP is broken, and $\epsilon$ controls the level of privacy, where a smaller $\epsilon$ signifies stronger privacy (and more noise). We evaluate the performance with various privacy budgets. Fig.14(a) shows that stronger privacy (a lower $\epsilon$) leads to lower final model accuracy, as the increased noise makes learning more difficult. Fig. 14(b) demonstrates that while all settings eventually reach a stable accuracy, stronger privacy constraints slow down the learning process. Fig. 14(c) illustrates that energy consumption is not significantly affected by the DP level. Overall, T-ELLM maintains robust performance when differential privacy is applied.
(a) Energy and accuracy
(b) Convergence of Wireless FL
(c) Energy consumption per round
Figure 13: Performance of the T-ELLM algorithm for wireless FL using gradient quantization.
(a) Energy and accuracy
(b) Convergence of Wireless FL
(c) Energy consumption per round
Figure 14: Performance of the T-ELLM algorithm for wireless FL using differential privacy.
VI-F Scalability for Large-Scale Scenario
(a) Energy efficiency
(b) Convergence of Wireless FL
(c) Energy consumption per round
Figure 15: Performance comparison of learning-based algorithms migrated to the scenario with $100$ devices.
To evaluate the scalability of T-ELLM, we simulate a larger weirless FL environment consisting of $100$ devices, where we distribute the CIFAR-10 dataset across these devices and configure the task to select $10$ participating devices in each training round. Meanwhile, we compare the performance of three algorithms:
  • T-ELLM Retrain: The T-ELLM agent is retrained using the virtual environment for the $100$-device scenario.
  • T-ELLM Original: The original T-ELLM agent, trained for the 20-device scenario, was directly applied without any retraining to test its zero-shot transferability.
  • SAC-Transformer: A baseline using a Transformer architecture with a linear head, trained from scratch on the $100$-device scenario using the SAC algorithm.
The results of these experiments are presented in Fig. 15. As shown in Fig. 15(a), the T-ELLM Retrain model achieves the highest energy efficiency and CWPEM value. Meanwhile, the T-ELLM Original model also performs well, demonstrating strong transferability. The SAC-Transformer baseline is comparatively less efficient. Fig.15(b) shows that both T-ELLM Retrain and T-ELLM Original achieve higher convergence accuracy than SAC-Transformer. In other words, the T-ELLM Original model can directly make effective device selections in a much larger environment. Finally, Fig.15(c) illustrates that the T-ELLM Retrain is the most energy-efficient. On the contrary, the SAC-Transformer consumes drastically more energy. These results indicate that the T-ELLM framework provides a robust and scalable solution for large-scale FL resource management.
VII Conclusion and Future Works
@@
We have presented T-ELLM, a novel tool-aided evolutionary LLM framework for efficient device selection and resource allocation in wireless FL. Notably, T-ELLM has been developed on top of a mathematics-driven decoupling of LLM-based device selection and tool-based resource allocation. Meanwhile, the combination of the linguistic reasoning capabilities in LLMs and mathematical optimization tools contributes to boosting the generalization capability of decision-making in environmental changes. In addition, T-ELLM takes advantage of a model-based virtual environment to support GRPO-based fine-tuning at minimal communication cost during interactions. Our theoretical analysis has proved the bounded discrepancy between virtual and real environments, which ensures the advantage function learned in the virtual environment maintains a provably small deviation from real-world conditions. Extensive experimental results have demonstrated that T-ELLM can further improve the energy efficiency and exhibit robust adaptability to environmental changes. Future work will focus on extending the T-ELLM framework to more complicated wireless resource management tasks, such as task offloading [47, 48]. Key priorities also include investigating its scalability, enhancing its safety against attacks, and validating its performance on real-world hardware. We will also continue to refine its algorithmic complexity to ensure practical deployment.